OpenStudy (anonymous):

prove that an nxn nilpotent matrix always has det=0

OpenStudy (anonymous):

I know that at some point A^n=0, and therefore the determinant is zero, but I must also have to prove that the matrices before it becomes zero have a determinant of zero

OpenStudy (anonymous):

and det(0)=0=A^n....and A^n is connected to A somehow

OpenStudy (anonymous):

use the fact det[A^n] = [det A]^n

OpenStudy (anonymous):

ok, so that would conclude the proof, because only 0^n can give zero

OpenStudy (anonymous):

just like x^n = 0, the only solution is x must be 0

OpenStudy (anonymous):

k thanks!